OpenStudy (anonymous):

Solve the system of equations 2x+4y=16 2x-4y=0

OpenStudy (lgbasallote):

add the equations..you end up with 4x = 16 divide both sides by 4 x = 4 got that?

OpenStudy (lgbasallote):

now sub in 4 into an x of any of the two equations then solve for y

OpenStudy (anonymous):

so is it (4,2) as an ordered pair

OpenStudy (lgbasallote):

no...like i said sub 4 into an x of any equation..let's say first equation...i'll substitute 4 into the x of the first equation 2y + 4x = 16 2y + 4(4) = 16 2y + 16 = 16 are you seeing the value of y now?

jimthompson5910 (jim_thompson5910):

you're right isabellayoung98, it's (4,2)

OpenStudy (lgbasallote):

oh lol..i misread sorry bout that

OpenStudy (anonymous):

thats wat i thought

OpenStudy (lgbasallote):

darn i switched y and x...i hate these bad eyes =_=

OpenStudy (lgbasallote):

im really sorry

OpenStudy (anonymous):

lol its okaii